Sol Summary – February 10th

Crew 221 Sol Summary Report 10Feb-2020


Summary Title: EVAs and Kim’s Release

Author’s name: Maria Grulich, M.Sc.

Mission Status: Nominal

Sol Activity Summary:

Update since last Sol summary: The crew had an astronomy lesson last night by crew astronomer G. Alotaibi, so all other crew members had a chance to ask questions and learn about it which was great!

This morning Crew Engineer L. Monge and Crew Commander M. Grulich went up the Hab Ridge road towards Hab Ridge North to get eight more microbiology samples as well as investigate how two different EVA teams with the same route and the same objective will perform.

The morning EVA already struggled to find the Hab Ridge road but managed after two extra little hill hikes. The team found great samples of stones with letch, several stratospheric layers and gryphaea at the top of Hab Ridge Road North, supporting the theory that Mars once hosted liquid water. We already know there are microbes living in the soil from the already incubated samples showing bacteria and fungi growth. Meanwhile, H.Kalucha and J. Todd worked on the robot and almost finished it, which just leaves the programming to be done tomorrow.

In the lunch break between the two EVA’s, we found out that we actually caught our little Martian and we called the rodent Kim and got permission to release him at Cowboys Corner at the end of EVA 9.

The second EVA of the day also had an issue finding the road, but still managed to find it and also collected 8 samples which will be processed today. Crew Commander prepared nutrient agar during the second EVA while being support lead to finish all samples by today to leave enough time for the incubation.

Once the second EVA team came back, they took the rovers to release Kim. The dinner was Rajma and Jalapeño Poppers by crew journalist H. Kalucha in preparation for Canadian/Indian Cultural night.

Look Ahead Plan: Planning of Operation EVAs, last microbiology sample preparation for the incubation and Canadian Cultural night.

Anomalies in work: None

Update mouse in hab: Caught Kim. We will set up the trap again to see if there is another one.

Weather: sunny partially cloudy in the afternoon light wind in the morning stronger wind in the afternoon

Crew Physical Status: Crew team members with signs of flu doing better, crew member with knee issues worsened after EVA

Reports to be filed:

Sol Summary

Photo Report

Operations Report

GreenHab Report

Science Report

Astronomy Report

EVA Request

EVA Report 8 and 9

Journalist Report

Support Requested: Raspberries same as yesterday Olive Oil from last time and White Vinegar if possible

Copyright © The Mars Society. All rights reserved. | Main Site